# Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice

### v1

@@ Abstract
Kinetic Mechanism for Itinerant Ferromagnetism.— To understand the underlying mechanism for the itinerant ferromagnetism, we examine the distinct microscopic hopping processes encompassed in the hopping term $H_{t}$ of the intermediate-$U$ Hubbard model, as schematically illustrated in Figs. 2(a)-(c). Although all these hopping processes are allowed, whether ferromagnetism is induced specifically by a single process or several cooperative processes has yet to be conclusively established. We introduce the projection operator $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and $\hat{P}_{\mathbf{i},{\sigma}}\equiv 1-\hat{Q}_{\mathbf{i},{\sigma}}$ to decompose the hopping term $H_{t}$:
@@
Consequently, $H_{t}$ includes NN hopping between (i) a doubly occupied site and an empty site with hopping amplitude $t_{\mathrm{PQ}}$, or two singly occupied sites with hopping amplitude $t_{\mathrm{QP}}$ [see Fig. 2(a)]; (ii) a singly occupied site and an empty site with hopping amplitude $t_{\mathrm{PP}}$ [see Fig. 2(b)], i.e., singlon hopping; (iii) a doubly occupied and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 2(c)], i.e., the doublon-singlon exchange. Given that case (i) describes Hermitian conjugate processes for an isotropic case, we consider $t_{\mathrm{PQ}}=t_{\mathrm{QP}}$. The hopping term in the Hubbard model (S1) corresponds to the isotropic limit $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=t_{\mathrm{QQ}}=t$. This decomposition enables us to examine the individual contributions of each microscopic hopping to itinerant ferromagnetism.
As shown in Figs. 2(e-f) for $L_{y}=4,6$ DMRG cylinders, before entering the fully polarized region where $m_{s}^{2}(\bm{\Gamma})$ saturates, multiple hopping processes have nonzero expectation values, suggesting that all contribute to the ground-state kinetic energy. In contrast, when increasing $U/t$, only $\langle H_{\mathrm{QQ}}\rangle$ is nonzero in the fully polarized region. This observation strongly suggests that, among distinct hopping processes, the doublon-singlon exchange uniquely drives the fully polarized ferromagnetism. Then we focus on this hopping process and further analyze the role of lattice geometric frustration in resulting ferromagnetism.
The effective NN spin interactions $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ stems from two kinds of virtual hopping processes at finite electron doping: (i) the superexchange $t^{2}/U$ favoring antiferromagnetic order, (ii) the doublon-assisted ring-exchange $t^{3}/V^{2}$ [see Fig. 2(d)] favoring either antiferromagnetic or ferromagnetic order depending on the sign of hopping $t$ Thouless (1965); Tasaki (1998a), where $V$ refers to the on-site potential pinning the doublon. Therefore, the effective spin interaction is $\sim\lim_{V\to 0}(t^{2}/U+{t^{3}}/{V^{2}})\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. This formulation reveals a fundamental difference between hole and electron doping in triangular lattices. For hole-doping ($t>0$), both hopping processes favor antiferromagnetism; while for electron-doping ($t<0$), the ring-exchange virtual process promotes ferromagnetism and finally dominates over the superexchange coupling in the large and infinite $U$ limit. Notably, square lattices lack such particle-hole asymmetry due to the positive ring-exchange coupling $\sim t^{4}$. This distinction highlights the important properties of geometrically frustrated lattices like the triangular lattice, where kinetic energy and geometric frustration work synergistically to foster itinerant ferromagnetism.
@@
Fig. 3: Impact of the microscopic hopping processes on itinerant ferromagnetism in the Hubbard model. (a)-(f) Static spin structure factor $S(\mathbf{q})$ ((a)-(c)) and the momentum distribution $n(\mathbf{k})$ ((d)-(f)) obtained by independently turning off each hopping process: (a),(d) $t_{\mathrm{PQ}}=0$, (b),(e), $t_{\mathrm{PP}}=0$, (c),(f), $t_{\mathrm{QQ}}=0$. Other hopping terms in (3) are set to $t$. The white hexagon indicates the first Brillouin zone; interpolation is applied to contours.
@@
To further verify the unique role of the doublon-singlon exchange in driving ferromagnetism, we independently deactivate each hopping amplitude while keeping the others unchanged, and compare the resulting ground states with the standard Hubbard model (S1). Therefore, the specific hopping process whose deactivation eliminates itinerant ferromagnetism is crucial for ferromagnetism. Figures 3 reveal that only switching off $t_{\mathrm{PQ}}$ or $t_{\mathrm{PP}}$ does not significantly affect the ferromagnetism, as the resulting $S(\mathbf{q})$ [see Figs. 3(a),(b)] and $n(\mathbf{k})$ [see Figs. 3(d),(e)] remain similar to those of the Hubbard model (S1) [see Figs. 1(c),(e)]. Conversely, switching off $t_{\mathrm{QQ}}$ eliminates ferromagnetic signatures [Figs. 3(c),(f)], since both the characteristic peak in $S(\mathbf{q})$ and well-defined Fermi surface in $n(\mathbf{k})$ disappear. These numerical results suggest the microscopic hopping process $H_{\mathrm{QQ}}$, as illustrated in Fig. 2(c), is pivotal in driving ferromagnetism at finite $U$. We further confirm this conclusion by examining a $t_{\mathrm{QQ}}-U$ model with only $t_{\mathrm{QQ}}=t$ and finite $U$ while $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=0$, and the itinerant ferromagnetism is observed (details see SupplementarySM ).
Evolution of Magnetism with Doping.— Below, we explore magnetism with a broader doping range. We examine both the local and global spin correlations through the NN spin correlations and the static spin structure factors, respectively.
@@
Fig. 4: Phase diagram. (a) Average NN spin correlations $\bar{S}(r=1)$ as a function of doping $\delta$ on width $L_{y}=4,6$ cylinders, resembling cold-atom quantum simulator findings Xu et al. (2023) for the same model. (b) Renormalized ferromagnetic order parameter $m_{s}^{2}(\mathbf{\Gamma})/m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})$ for the Hubbard model (S1) as a function of doping $\delta$ and interaction $U/t$. $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$ is full polarization value of $m_{s}^{2}(\mathbf{\Gamma})$. Circles denote the data obtained numerically.
@@
The average local spin correlations are defined as $\bar{S}(r)=\frac{1}{z}\sum_{\mathbf{j},|\mathbf{i}_{0}-\mathbf{j}|=r}\langle\mathbf{S}_{\mathbf{i}_{0}}\cdot\mathbf{S}_{\mathbf{j}}\rangle,$ with the summation over neighbors at a distance $r$ from a fixed site $\mathbf{i}_{0}$ that is strategically positioned at the lattice’s center to avoid a boundary effect. Here, we consider $r=1$ for NN spin correlations, and $z=6$ represents the coordination number of the triangular lattice. As shown in Fig. 4(a), the average NN correlations are antiferromagnetic across a broad range of hole doping ($\delta<0$) at intermediate coupling strength $U/t=10$, yet they are markedly suppressed by electron doping ($\delta>0$). In particular, we observe that the local spin correlations become ferromagnetic at intermediate electron doping. Notably, these local spin correlations at zero temperature align well with the recent cold-atom experimental findings (Fig. 3b of Ref. Xu et al. (2023)) at finite temperature for the similar intermediate $U/t$. This observation suggests that the local magnetic properties of the ground state survive in a finite range of temperatures.
@@
We further examine the static spin structure factors to establish the magnetic phase diagram as a function of interaction $U/t$ and electron doping $\delta>0$. As illustrated in Fig.4(b) for $L_{y}=6$ cylinders, the renormalized squared order parameter $m_{s}^{2}(\bm{\Gamma})/m_{s,F}^{2}(\bm{\Gamma})$ shows ferromagnetism in a finite range of dopings $\delta\approx 1/2$ and interactions $U\gtrsim 10$, which smoothly extends to the infinite interaction limit. These findings are consistent with $L_{y}=4$ cylinders SM . Together with the numerical evidence presented in the previous section, this demonstrates that the ferromagnetism at both finite and infinite $U$ around $\delta\approx 1/2$ is driven by the doublon-singlon exchange.
@@
Moreover, for finite $U$ and dopings both below and above $\delta\approx 1/2$, the incommensurate spin density wave (iSDW) emerges SM for both $L_{y}=4$ and $L_{y}=6$, which separates the ferromagnetic phase at $n=3/2$ from the phases near the full-filling and half-filling limits. We remark that, unlike the consistent observations at finite coupling strength, at $U=\infty$, we find a smooth connection between the ferromagnetic phase at $\delta\gtrsim 0$ and $\delta\approx 1/2$ when $L_{y}=6$, but these two phases are separated by an iSDW phase when $L_{y}=4$ SM . In addition, for larger doping $15/8\lesssim\delta<$ 1, both $L_{y}=4$ and $L_{y}=6$ systems exhibit another iSDW phase [see supplementarySM ], separating the ferromagnetic regions observed at $\delta\approx 1/2$ from ferromagnetism at $\delta\to 1$, thus suggesting distinct underlying mechanisms in these two doping concentrations.
Summary and outlook.— In this work, we explore the ferromagnetism mechanism of the finite-doped triangular-lattice Hubbard model at intermediate on-site Coulomb interaction. By diagnosing each microscopic hopping process separately, we reveal that the geometrically frustrated doublon-singlon exchange among other hopping processes in triangular lattices solely contributes to the fully polarized ferromagnetism. We further establish the magnetic phase diagram for the electron-doped triangular Hubbard model and compare the local spin correlations with the latest cold-atom experiments. Our findings may stimulate future studies of the triangular-lattice ferromagnetism beyond the cold-atom systems, particularly in condensed matter physics. For instance, the TMD moiré materials offer platforms with additional tuning parameters to study the stability of ferromagnetism against long-range Coulomb interactions in TMD hetero-bilayers Jin et al. (2021); Xu et al. (2020); Senthil et al. (2004a, b); Senthil (2023) and complex hopping in TMD homo-bilayers Wang et al. (2020). It may also inspire future finite-temperature numerical works to understand the relationship between rising temperature and increasing coupling strength, as previously focused on the half-filling case Wietek et al. (2021); Zhang et al. (2022).
Acknowledgements.
We particularly thank helpful discussions with Ashvin Vishwanath. This work is supported by the National Natural Science Foundation of China (Grant No.12074375), the Fundamental Research Funds for the Central Universities and the Strategic Priority Research Program of CAS (Grant No.XDB33000000).
Supplementary Materials for
“Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice”
I DMRG Details and DMRG Convergence

### v2

@@ RESULTS
Microscopic mechanism: geometrical frustration assisted kinetic ferromagnetism. In the above, we have identified the emergence of itinerant ferromagnetism with increasing $U/t$ at $\delta=1/2$. To understand the underlying mechanism for the itinerant ferromagnetism, we examine the distinct microscopic hopping processes encompassed in the hopping term $H_{t}$ of the intermediate-$U$ Hubbard model, as schematically illustrated in Figs. 3a, b, c. We introduce the projection operator $\hat{Q}_{\mathbf{i},{\sigma}}\equiv n_{\mathbf{i},{\sigma}}$ and $\hat{P}_{\mathbf{i},{\sigma}}\equiv 1-\hat{Q}_{\mathbf{i},{\sigma}}$ to decompose the hopping term $H_{t}$:
@@
Consequently, $H_{t}$ includes NN hopping between (i) a doubly occupied site and an empty site with hopping amplitude $t_{\mathrm{PQ}}$, or two singly occupied sites with hopping amplitude $t_{\mathrm{QP}}$ [see Fig. 3a]; (ii) a singly occupied site and an empty site with hopping amplitude $t_{\mathrm{PP}}$ [see Fig. 3b], i.e., singlon hopping; (iii) a doubly occupied and a singly occupied site with hopping amplitude $t_{\mathrm{QQ}}$ [see Fig. 3c], i.e., the doublon-singlon exchange. Given that case (i) describes Hermitian conjugate processes for an isotropic case, we consider $t_{\mathrm{PQ}}=t_{\mathrm{QP}}$. The hopping term in the Hubbard model (1) corresponds to the isotropic limit $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=t_{\mathrm{QQ}}=t$. In contrast to the large $U$ limit, where the $t$-$J$ model effectively describes the Hubbard model with only the $t_{\mathrm{QQ}}$ hopping term in the no-empty-occupancy Hilbert space for electron doping ($n>1$) ShuaiChen2022SCBA , intermediate $U$ regimes are generally regarded as involving multiple kinetic processes that cannot be reduced to the $t$-$J$ model alone Spalek2007 ; KeYang2024 . However, it remains unclear whether ferromagnetism at intermediate $U$ is induced by a single kinetic process or by the cooperative effect of several processes.
The decomposition (6) enables us to examine the individual contributions of each microscopic hopping to itinerant ferromagnetism. As shown in Figs. 3e, f for $L_{y}=4,6$ DMRG cylinders, before entering the fully polarized region where $m_{s}^{2}(\boldsymbol{\Gamma})$ saturates, multiple hopping processes have nonzero expectation values, suggesting that they contribute to the ground-state kinetic energy. In contrast, when increasing $U/t$ to intermediate coupling, only $\langle H_{\mathrm{QQ}}\rangle$ is nonzero in the fully polarized region. We remark that, unlike the infinite $U$ or strong-coupling limits, the result in Figs. 3e, f is particularly unexpected for intermediate coupling ($U/t\sim 10$) where multiple kinetic processes are typically significant. These observations illustratively suggests that, among distinct hopping processes in the decomposition (6), the doublon-singlon exchange dominantly contribute to the kinetic energy of fully polarized ferromagnetism. While this is conceptually similar to Nagaoka’s ferromagnetism, the conditions for the emergence of the ferromagnetism—intermediate interaction ($U\gtrsim 10$) and intermediate doping ($\delta=1/2$)—significantly violate the strict preconditions of Nagaoka’s theorem. In order to compare to the standard Nagaoka regime (single electron doping and $U=\infty$), we repeated the Fig. 3e analysis for a single electron doped $8\times 4$ cylinder and found that, even at $U/t=260$, the ground state remains only partially polarized. Nonetheless, similar behavior persists: both $\langle H_{\mathrm{PQ}}+H_{\mathrm{QP}}\rangle$ and $\langle H_{\mathrm{QQ}}\rangle$ contribute at intermediate couplings, while the doublon-singlon exchange term $\langle H_{\mathrm{QQ}}\rangle$ steadily dominates as $U/t$ increases. This can be naturally extrapolated to the $U\to\infty$ limit, where the doublon-singlon exchange $\langle H_{\mathrm{QQ}}\rangle$ becomes the sole active kinetic process in the fully polarized ferromagnetic state. We then focus on doublon-singlon exchange and show that the doublon-assisted ring-exchange [see Fig. 3d] leads to the ferromagnetism, unlike the antiferromagnetism typically driven by superexchange.
@@
Fig. 4: $|$ Impact of the microscopic hopping processes on itinerant ferromagnetism in the Hubbard model. (a)-(f) Static spin structure factor $S(\mathbf{q})$ ((a)-(c)) and the momentum distribution $n(\mathbf{k})$ ((d)-(f)) obtained by independently turning off each hopping process: (a),(d) $t_{\mathrm{PQ}}=0$, (b),(e), $t_{\mathrm{PP}}=0$, (c),(f), $t_{\mathrm{QQ}}=0$. Other hopping terms in Eq. (6) are set to $t$. The white hexagon indicates the first Brillouin zone; interpolation is applied to contours. Here, we fix the system size as $8\times 4$ cylinders, the electron doping $\delta=1/2$, and the interaction strength $U=10$.
The effective NN spin interactions stem from two kinds of virtual hopping processes at finite electron doping: (i) the superexchange $t^{2}/U$ favoring antiferromagnetic order, (ii) the doublon-assisted ring-exchange $t^{3}/\Delta E^{2}$ favoring either antiferromagnetic or ferromagnetic order depending on the sign of hopping $t$ Thouless1965 ; Tasaki1998 , where $\Delta E$ refers to the energy difference between the initial and intermediate states. Therefore, the effective spin interaction is $\sim(t^{2}/U+{t^{3}}/{\Delta E^{2}})\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. This formulation reveals a fundamental difference between hole and electron doping in triangular lattices. For hole-doping ($t>0$), both hopping processes favor antiferromagnetism; while for electron-doping ($t<0$), the ring-exchange virtual process promotes ferromagnetism and finally dominates over the superexchange coupling in the large and infinite $U$ limit. Notably, square lattices lack such particle-hole asymmetry due to the positive ring-exchange coupling $\sim t^{4}$. This distinction highlights the important properties of geometrically frustrated lattices like the triangular lattice, where kinetic energy and geometric frustration work synergistically to foster itinerant ferromagnetism. In addition to the above perturbation analysis, we also provide an alternative explanation grounded in the symmetry of wave functions to understand the ferromagnetism origin (see Supplementary Note 2).
@@
To show stronger evidence that the doublon-singlon exchange process drives ferromagnetism, we independently deactivate each hopping amplitude while keeping the others unchanged, and compare the resulting ground states with the standard Hubbard model (1). In this manner, the specific hopping process whose deactivation eliminates itinerant ferromagnetism is decisive for ferromagnetism. Figures 4 reveal that only switching off $t_{\mathrm{PQ}}$ or $t_{\mathrm{PP}}$ does not significantly affect the ferromagnetism, as the resulting $S(\mathbf{q})$ [see Figs. 4a, b] and $n(\mathbf{k})$ [see Figs. 4d, e] remain similar to those of the Hubbard model (1) [see Fig. 1e and Fig. 2b]. Conversely, switching off $t_{\mathrm{QQ}}$ eliminates ferromagnetic signatures [Figs. 4c, f], since both the characteristic peak in $S(\mathbf{q})$ and well-defined Fermi surface in $n(\mathbf{k})$ disappear. These numerical results suggest the microscopic hopping process $H_{\mathrm{QQ}}$, as illustrated in Fig. 3c, is necessary in driving ferromagnetism at finite $U$. We further confirm this conclusion by examining a $t_{\mathrm{QQ}}-U$ model, where only $t_{\mathrm{QQ}}=t$ and finite $U$ are retained, while $t_{\mathrm{PQ}}=t_{\mathrm{QP}}=t_{\mathrm{PP}}=0$. In this scenario, the itinerant ferromagnetism is observed (see details in Supplementary Note 3).
@@
Fig. 5: $|$ Magnetic phase diagram and stability. (a) Renormalized ferromagnetic order parameter $m_{s}^{2}(\mathbf{\Gamma})/m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})$ for the Hubbard model (1) as a function of electron doping $\delta$ and $U/t$, where $U$ is the on-site Hubbard interaction, and $t$ is the nearest-neighbor (NN) hopping amplitude. Here, $m_{s,\mathrm{F}}^{2}(\mathbf{\Gamma})\equiv[(1-\delta)/2]^{2}$ is full polarization value of $m_{s}^{2}(\mathbf{\Gamma})$. Circles denote the data obtained numerically. (b) Average NN spin correlations $\bar{S}(r=1)$ as a function of doping $\delta$ on cylinders of width $L_{y}=4,6$, resembling cold-atom quantum simulator findings MuqingXu2023 for the same model. (c) Static structure factor $S(\mathbf{q})$ at the $\boldsymbol{\Gamma}$ point as a function of $J_{3}$, for the Hubbard model with additional term $\delta H=-J_{3}\sum_{\langle\mathbf{i}\mathbf{j}\rangle}S_{\mathbf{i}}^{z}S_{\text{$\mathbf{j}$}}^{z}$ at $U=10$ and $8\times 4$ cylinder. The ferromagnetism persists for $J_{3}>0$. Inset: $S(\mathbf{q})$ for $J_{3}=3$.
Magnetic Phase Diagram. Now we explore magnetism with a broader doping range and examine both the local and global spin correlations.
@@
The average local spin correlations are defined as $\bar{S}(r)=\frac{1}{z}\sum_{\mathbf{j},|\mathbf{i}_{0}-\mathbf{j}|=r}\langle\mathbf{S}_{\mathbf{i}_{0}}\cdot\mathbf{S}_{\mathbf{j}}\rangle,$ with the summation over neighbors at a distance $r$ from a fixed site $\mathbf{i}_{0}$ that is strategically positioned at the lattice’s center to avoid a boundary effect. Here, we consider $r=1$ for NN spin correlations, and $z=6$ represents the coordination number of the triangular lattice. As shown in Fig. 5b, the average NN correlations are antiferromagnetic across a broad range of hole doping ($\delta<0$) at intermediate coupling strength $U/t=10$, yet they are markedly suppressed by electron doping ($\delta>0$). In particular, we observe that the local spin correlations become ferromagnetic at intermediate electron doping. Notably, these local spin correlations at zero temperature align well with the recent cold-atom experimental findings (figure 3b of Ref. (MuqingXu2023, )) at finite temperature for the similar intermediate $U/t$. This observation suggests that the local magnetic properties of the ground state survive in a finite range of temperatures.
@@
We further examine the static spin structure factors to establish the magnetic phase diagram as a function of interaction $U/t$ and electron doping $\delta>0$. As illustrated in Fig. 5a for $L_{y}=6$ cylinders, the renormalized squared order parameter $m_{s}^{2}(\boldsymbol{\Gamma})/m_{s,F}^{2}(\boldsymbol{\Gamma})$ shows ferromagnetism in a finite range of dopings $\delta\approx 1/2$ and interactions $U\gtrsim 10$, which smoothly extends to the infinite interaction limit. These findings are consistent with $L_{y}=4$ cylinders (see Supplementary Note 4). Together with the numerical evidence presented in the previous section, this demonstrates that the ferromagnetism at both finite and infinite $U$ around $\delta\approx 1/2$ is driven by the doublon-singlon exchange.
@@
Moreover, for finite $U$ and electron doping both below and above $\delta\approx 1/2$, the iSDW emerges (see Supplementary Note 4) for both $L_{y}=4$ and $L_{y}=6$, which separates the ferromagnetic phase at $n=3/2$ from the phases near the full-filling and half-filling limits. We remark that, unlike the consistent observations at finite coupling strength, at $U=\infty$, we find a smooth connection between the ferromagnetic phase at $\delta\gtrsim 0$ and $\delta\approx 1/2$ when $L_{y}=6$, but these two phases are separated by an iSDW phase when $L_{y}=4$ (see Supplementary Note 4). In addition, for larger doping, both $L_{y}=4$ and $L_{y}=6$ systems exhibit another iSDW phase (see Supplementary Note 4), separating the ferromagnetic regions observed at $\delta\approx 1/2$ from ferromagnetism at $\delta\to 1$, thus suggesting distinct underlying mechanisms in these two doping concentrations.
Stable Ferromagnetism: Magnetic Anisotropy. The ferromagnetic ground state of the Hubbard model spontaneously breaks SU(2) symmetry. Given that stable ferromagnetism is crucial for the development of spintronic devices based on high-temperature ferromagnetic materials, as reported in recent experiments, we create a spin-wave excitation gap by introducing magnetic anisotropy to stabilize this state. Specifically, we consider the following anisotropic term added to the Hubbard Hamiltonian $\delta H=-J_{3}\sum_{\langle\mathbf{i}\mathbf{j}\rangle}S_{\mathbf{i}}^{z}S_{\text{$\mathbf{j}$}}^{z}$, which can be readily realized by experiments. Here we require $J_{3}>0$. Using the random phase approximation (RPA) on the Hartree-Fock ground state, we show that the anisotropic term $\delta H$ introduces a gap in the excitation spectrum (see details in Supplementary Note 5), which secures the ferromagnetism even at finite temperatures.
We illustrate the stability of FM phase based on the Hartree-Fock (HF) approximation. We start with the HF Hamiltonian for a FM state with magnetism $m$
$H_{\mathrm{HF}}$ $=\sum_{\mathbf{k},\sigma}(\epsilon_{\mathbf{k}}-\mu)c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma}-\frac{1}{2}Um\sum_{\mathbf{k}}c_{\mathbf{k},\alpha}^{\dagger}{\sigma_{\alpha\beta}^{z}}c_{\mathbf{k},\beta}$
$=\sum_{\mathbf{k},\sigma}(\epsilon_{\mathbf{k}}-\mu-\frac{1}{2}Um\sigma)c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma}$
$\equiv\sum_{\mathbf{k}\sigma}\xi_{\mathbf{k},\sigma}c_{\mathbf{k},\sigma}^{\dagger}c_{\mathbf{k},\sigma},$ (7)
where $\sigma=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the Pauli matrix. We denote the FM ground state of the HF Hamiltonian as $|\Omega\rangle$. Using the random phase approximation (RPA) on the Hartree-Fock ground state, we could derive the spin susceptibility
$\chi_{\mathrm{RPA}}^{ij}(\mathbf{q},\omega)=-\frac{\chi_{0}^{+-}(\mathbf{q},\omega)}{\frac{\omega}{mU}-\frac{q^{2}}{2m_{e}mU}}\delta_{ij},\quad i,j=x,y,$ (8)
with $m_{e}$ being the effective mass of electrons. Here $\chi_{0}^{+-}(\mathbf{q},\omega)$ is the bare spin susceptibility at $xx$, yyyy directions [see Supplementary Note 5]. The susceptibility indicates a spin wave of the square dispersion $\omega(\mathbf{q})\propto q^{2}$. We can decompose the perturbation term $\delta H$, which gives rise to the extra contribution $\delta H_{\mathrm{HF}}$ to the Hartree-Fock Hamiltonian
$\delta H_{\mathrm{HF}}=-2zmJ_{3}\sum_{\mathbf{i}}S_{\mathbf{i}}^{z}+zm^{2}J_{3}.$ (9)
Here, $z$ is the coordinate number. Then the RPA spin susceptibility accordingly changes to
$\chi_{\mathrm{RPA}}^{ij}(\mathbf{q},\omega)=-\frac{\chi_{0}^{+-}(\mathbf{q},\omega)}{\frac{\omega}{mU}-\frac{q^{2}/(2m_{e})+2zJ_{3}}{mU}}\delta_{ij},\quad i,j=x,y.$ (10)
Obviously, a finite gap $\Delta_{\mathrm{gap}}=2zJ_{3}$ is introduced in the spin waves, which can suppress the thermal fluctuations of the spin wave as such to stabilize a FM at finite temperature. We further validate the stability of ferromagnetism under the influence of the anisotropic term $\delta H$ through DMRG simulations across a range of $J_{3}$ values. As illustrated in Fig. 5c, we analyze the static structure factor at the $\boldsymbol{\Gamma}$ point as a function of $J_{3}$. Our observations indicate that ferromagnetism remains robust for $J_{3}>0$. Together with the presence of the gap in the dispersion, the DMRG results suggest that ferromagnetism remains stable after the introduction of anisotropy.
Discussion
In this work, we investigate the ferromagnetism mechanism of the finite-doped triangular-lattice Hubbard model with intermediate on-site Coulomb interaction, identifying the cooperative interplay between the charge kinetic process and lattice geometry as the driving factors of ferromagnetism. By separately diagnosing each of the four microscopic hopping processes decomposed from Hubbard model, we reveal that the geometrically frustrated doublon-singlon exchange is the only hopping process in triangular lattices for achieving the fully polarized ferromagnetism. We also establish the magnetic phase diagram for the electron-doped triangular Hubbard model and compare the local spin correlations with the latest experiments. urthermore, we demonstrate that the ferromagnetic phase remains robust even when explicit $SU(2)$ symmetry breaking is introduced via anisotropy.
Our findings may serve as the starting point and stimulate future studies closely related to distinct TMD hetero- and homo-bilayer materials, such as exploring the stability of ferromagnetism against long-range Coulomb interactions and complex hoppings. Notably, an insulating stripe phase at the $\delta=1/2$ electron doping was recently found experimentally on transition-metal dichalcogenides (TMD) moiré superlattices formed by hetero-bilayers of WSe2/WS2 JinChenhao2021 , which are believed to be described by the extended Hubbard models in the strong coupling limit with long-range Coulomb interactions $V$. This revelation invites further inquiries into the stability of ferromagnetism under the influence of long-range Coulomb interactions, the emergence of other correlated phases XuYang2020 , and the possible direct quantum phase transitions among these conventional symmetry-breaking phases Senthil2004Science ; Senthil2004 ; Senthil2023 . In addition, another type of TMD moiré superlattices formed by homo-bilayers LeiWang2020 has been proposed to be described by the moiré Hubbard model, which generalizes the pure Hubbard model by introducing complex phases in hoppings $te^{i\sigma\phi}$. Therefore, our work may further motivate the investigations into how the hopping phase influences the emergence of kinetic-energy-driven ferromagnetism. Moreover, these findings may also shed light on the magnetism with general interaction in diverse lattice geometries Mielke1991 ; Mielke1992 ; Wurth1996 ; ChiaChenChang2010 ; Scholle2023 ; Riegler2023 ; Scholle2024 .